rell:

1. 5( 3x + 2 ) 2. 4( a - b ) 3. -3( 2r2 + 7s ) 4. 2x( 5x - 6 ) can yall give me the answer to these thnk u

snowflake0531:

do you know the distributive property?

rell:

no

ItzJasminefools:

whats 3+2?

snowflake0531:

so a(b+c) = ab+ac a(b-c) = ab-ac

ItzJasminefools:

oop

100kpimpcess:

oof~

snowflake0531:

by the way, do you know how to multiply variables and numbers together?

100kpimpcess:

number 1 u wanna get x by itself

100kpimpcess:

bc x is being multiplied by 3 which is really x (1)

carmelle:

we have to distribute first @100kpimpcess

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle we have to distribute first @100kpimpcess \(\color{#0cbb34}{\text{End of Quote}}\) ye

rell:

ik how to do it

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rell ik how to do it \(\color{#0cbb34}{\text{End of Quote}}\) ight u sure

100kpimpcess:

@darkknight

rell:

yes

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rell yes \(\color{#0cbb34}{\text{End of Quote}}\) ight

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 so a(b+c) = ab+ac a(b-c) = ab-ac \(\color{#0cbb34}{\text{End of Quote}}\) good example of distributive property, use this listed above you can distribute all of the above equations, try it @rell?